# latexpage numbers missing

### v1

@@ 5 Acknowledgments
Fix parameter $\bm{\theta}=(\lambda,\nu)$.
Select a minimum and maximum number of observed locations, $o_{L},o_{U}\in\mathbb{N}^{+}$.
for $o_{j}$ in $o_{L}:o_{U}$ do
$\{\bm{x}_{0,k}\}_{k\in[s]}\sim p(\cdot\mid\bm{\theta})$.
@@
  $\mathring{\mathcal{S}}_{l}=\{\bm{s}_{l_{i}}\in\mathcal{S}\mid\bm{s}_{l_{i}}\sim\textrm{Unif}(\mathcal{S}),i\in[o_{j}]\}$
  $t_{l}\sim U([0,T])\textrm{ for }l\in[m],\quad[0,T]\in\mathbb{Z}$
  $\bm{\mathring{x}}_{0,l}=\bm{M}(\mathring{\mathcal{S}}_{l})\odot\bm{x}_{0,k}\textrm{ for }l\in[m]$
  $\tilde{\bm{x}}_{t_{l}}\sim p(\cdot\mid\bm{\mathring{x}}_{0,l},\tilde{\bm{x}}_{0,k}):=\mathcal{N}(\sqrt{\tilde{\alpha}_{t}}\tilde{\bm{x}}_{0,k},\tilde{\sigma}_{t}^{2}\bm{I})\textrm{ for }l\in[m]$
  $\mathcal{D}\leftarrow\mathcal{D}\cup\{\big{(}(\bm{\mathring{x}}_{0,l},\tilde{\bm{x}}_{t_{l}}),\bm{M}(\mathring{\mathcal{S}}_{l}),\bm{\theta}_{j},t_{l}\big{)}\mid l\in[m]\}$
@@
S2.3 Validation Details Common to Both Case Studies
Conditional Evaluation Dataset (Proportion U-Net)
For each of the five observed proportions $\rho\in\{0.01,0.05,0.1,0.25,0.5\}$, we generated a conditioning set $\mathring{\mathcal{S}}$ and mask $\bm{M}(\mathring{\mathcal{S}})$ using the Bernoulli distribution with probability $\rho$ of observing each location and a reference spatial field $\bm{x}_{0}$ using the unconditional distribution with either variance $\nu=1.5$ and $\ell=3$ (Gaussian process) and smoothness $\nu=1.5$ and range $\lambda=3$ (Brown–Resnick process). For each of the resulting five partially observed spatial fields $\bm{\mathring{x}}_{0}=\bm{M}(\mathring{\mathcal{S}})\odot\bm{x}_{0}$, we generated $m=4000$ replicates of the unobserved spatial field $\{\tilde{\bm{x}}_{0,i}\}_{i\in[m]}$ with NCS and either the true predictive distribution or an MCMC approximation depending on the spatial process.
Conditional Evaluation Dataset (Parameter U-Net)
For each length scale
$\ell\in\{1,2,3,4,5\}$}, we generated a reference spatial field $\bm{x}_{0}$ using the Gaussian process unconditional simulator $p(\cdot\mid\bm{\theta})$ for the given length scale and fixed variance $\nu=1.5$. We repeated this process for the Brown–Resnick process with range values $\lambda\in\{1,2,3,4,5\}$ and fixed smoothness parameter $\nu=1.5$. For each reference spatial field, we generated a conditioning set $\mathring{\mathcal{S}}$ and corresponding mask $\bm{M}(\mathring{\mathcal{S}})$ using the Bernoulli distribution with probability $\rho=0.05$ of observing each location. Using each of the five partially observed fields $\bm{\mathring{x}}_{0}=\bm{M}(\mathring{\mathcal{S}})\odot\bm{x}_{0}$, we generated $m=4000$ unobserved spatial field replicates $\{\tilde{\bm{x}}_{0,i}\}_{i\in[m]}$ with NCS and either the true predictive distribution or an MCMC approximation depending on the spatial process.
Conditional Evaluation Dataset (Small Conditioning Set)
@@
For each range value $\lambda\in\{1,2,3,4,5\}$ and each conditioning set size $m$ ranging from one to seven, we generated a reference spatial field $\bm{x}_{0}$ using the Brown–Resnick process unconditional simulator $p(\cdot\mid\bm{\theta})$ for the given range $\lambda$ and fixed smoothness $\nu=1.5$. For each reference spatial field, we generated a conditioning set $\mathring{\mathcal{S}}$ and corresponding mask $\bm{M}(\mathring{\mathcal{S}})$ through discrete uniform sampling without replacement of $\mathcal{S}$ such that $\lvert\mathring{\mathcal{S}}\rvert=m$. Using each of the $35$ partially observed fields $\bm{\mathring{x}}_{0}=\bm{M}(\mathring{\mathcal{S}})\odot\bm{x}_{0}$, we generated $m=4000$ unobserved spatial field replicates $\{\tilde{\bm{x}}_{0,i}\}_{i\in[m]}$ with NCS and FCS—an MCMC approximation described in Section 3.1.
Unconditional Evaluation Dataset (Parameter and Proportion U-Net)
We apply the process described in Section 2.6 and shown in Algorithm S3 using either varying range $\lambda\in\{1,2,3,4,5\}$, fixed smoothness $\nu=1.5$, and fixed observed proportion $\rho=.05$ for the parameter U-Net or varying observed proportion $\rho\in\{0.01,0.05,0.1,0.25,0.5\}$ with fixed range $\lambda=3$ and smoothness $\nu=1.5$ for the proportion U-Net.
@@
Algorithm S3 Generating NCS-approximated Unconditional Evaluation Dataset (Parameter and Proportion U-Nets)
@@
Fix parameter $\bm{\theta}=(\lambda,\nu)$ and observed proportion $\rho$ and set $\mathcal{D}_{\textrm{eval}}=\emptyset$
for $i$ in $1:m$ do
@@
$\bm{M}(\mathring{\mathcal{S}}_{i})\sim B(n,\rho)$ with $n=32^{2}$.
$\tilde{\bm{x}}_{0,i}\sim\tilde{p}_{0}(\cdot\mid\bm{\mathring{x}}_{0,i},\bm{\theta})\textrm{ for }\bm{\mathring{x}}_{0,i}=\bm{M}(\mathring{\mathcal{S}}_{i,0})\odot\bm{x}_{0,i}\textrm{ and NCS/LCS distribution }\tilde{p}_{0}(\cdot\mid\bm{\mathring{x}}_{0,i},\bm{\theta})$
$\mathcal{D}_{\textrm{eval}}\leftarrow\mathcal{D}_{\textrm{eval}}\cup\{[\bm{\mathring{x}}_{0,i},\tilde{\bm{x}}_{0,i}]\}$
@@
Return $\mathcal{D}_{\textrm{eval}}=\{[\bm{\mathring{x}}_{0,i},\tilde{\bm{x}}_{0,i}]\}_{i\in[m]}$
Unconditional Evaluation Dataset (Small Conditioning Set U-Net)
We apply the process described in Section 2.6 and shown in Algorithm S4 using varying range $\lambda\in\{1,2,3,4,5\}$}, fixed smoothness $\nu=1.5$, and conditioning sets with varying sizes ranging from one to seven.
Algorithm S4 Generating NCS-approximated Unconditional Evaluation Dataset (Small Conditioning set)
Fix parameter $\bm{\theta}=(\lambda,\nu)$ and number of observed locations $o_{j}$ and set $\mathcal{D}_{\textrm{eval}}=\emptyset$
for $i$ in $1:m$ do
@@
$\mathring{\mathcal{S}}_{i}\sim\textrm{Unif}(\bm{\mathcal{S}}_{o_{j}})\textrm{ where }\bm{\mathcal{S}}_{o_{j}}=\{\mathcal{S}_{i}\subset\mathcal{S}\mid\lvert\mathcal{S}_{i}\rvert=o_{j}\}$
$\tilde{\bm{x}}_{0,i}\sim\tilde{p}_{0}(\cdot\mid\bm{\mathring{x}}_{0,i},\bm{\theta})\textrm{ for }\bm{\mathring{x}}_{0,i}=\bm{M}(\mathring{\mathcal{S}}_{i,0})\odot\bm{x}_{0,i}\textrm{ and NCS/FCS distribution }\tilde{p}_{0}(\cdot\mid\bm{\mathring{x}}_{0,i},\bm{\theta})$
$\mathcal{D}_{\textrm{eval}}\leftarrow\mathcal{D}_{\textrm{eval}}\cup\{[\bm{\mathring{x}}_{0,i},\tilde{\bm{x}}_{0,i}]\}$
@@
Return $\mathcal{D}_{\textrm{eval}}=\{[\bm{\mathring{x}}_{0,i},\tilde{\bm{x}}_{0,i}]\}_{i\in[m]}$
Conditional Univariate and Bivariate Densities
Conditional univariate and bivariate densities are computed using the conditional evaluation datasets and kernel density estimation with a Gaussian kernel and bandwidth selected according to Scott’s Rule (Scott, 1992). Specifically, we display a $3\times 5$ grid of plots in which the first row contains the partially observed spatial fields and the second and third rows contain the conditional univariate and bivariate kernel density estimates respectively for the selected unobserved locations. In the first row, red dots (univariate) and black triangles (bivariate) mark the selected spatial locations. The true observed values at the selected locations are marked with red, dashed vertical lines. From left to right along each column, either the parameter of interest or the observed proportion is increasing.
Conditional Mean Fields
The empirical conditional mean fields are computed by averaging the $m=4000$ replicates at each of the unobserved locations in the conditional evaluation dataset. We display a $3\times 5$ grid of plots that contains the partially observed spatial fields in the first row, the true empirical (Gaussian) or an MCMC-approximated (Brown–Resnick) conditional mean field in the second row, and the empirical NCS conditional mean field in the third row. In the bottom two rows, the observed locations in the partially observed field used to generate the empirical conditional mean fields are masked.
Conditional Correlation Metrics
To evaluate how well the NCS simulations capture the spatial correlations of the true predictive distributions, we produce empirical conditional correlation heatmaps for the Gaussian process case study. For a fixed unobserved spatial location in a partially observed spatial field from the conditional evaluation dataset, we compute the Pearson correlation coefficient (PCC) between the unobserved spatial location and all other unobserved spatial locations using either the NCS or true conditional simulations. PCC measures the linear correlation between two unobserved spatial locations $\bm{s}_{1},\bm{s}_{2}$ given the partially observed spatial field. If the PCC is 1-1 or $1$, there is perfect linear correlation given the partially observed spatial field whereas a PCC value of 0 indicates no linear correlation given the partially observed field. Specifically, we display a $2\times 5$ grid of plots in which the first row contains the true empirical correlation heat maps and the second row contains the equivalent for NCS.
@@
The previous metrics such as the conditional univariate and bivariate densities, conditional mean fields, and conditional correlation metrics only reveal how the low-dimensional properties of the NCS and the true predictive distributions or standard approximations compare. Yet, we can implicitly evaluate whether the NCS predictive distributions exhibit similar high-dimensional properties as the true predictive distributions even if the true predictive distributions are intractable using the unconditional evaluation dataset and multivariate metrics. For the Brown–Resnick case study, we examined a variety of multivariate metrics yet only display a select few in this paper. The extremal correlation function and distributions of the spatial minimum, maximum, and absolute summation of all values in each full spatial field replicate from the unconditional evaluation dataset are shown in a $4\times 3$ grid of plots.
S3 Gaussian Process: Tractable Case Study
We applied NCS to a Gaussian process with a zero mean function and an exponential covariance function with two positive parameters—the length scale $\ell$ and the variance $\nu$. Hence, the model is
[…]
S3.1 Training Details
From experimentation, there are few hyperparameters that are sensitive to how well the U-Net approximates the true conditional score function. As such, the reason the hyperparameters vary between the U-Net and spatial process types is due to computational constraints, not training issues. The only training sensitivity we encountered pertains to amortization—the amortized variable ought to be sampled from a continuous, positive distribution.
Parameter U-Net
The variance is fixed at $\nu=1.5$, the length scale $\ell\in[1.0,5.0]$ varies, and the proportion of observed locations $\rho=0.05$ is fixed. To simulate training and validation data, we follow the process described in Section S2.2. The training data has the following parameters: $\rho=0.05$, $p=1000$ parameters samples from $\tilde{\Theta}=[0.5,5.5]\times\{1.5\}$, $s$ = 1 full spatial field realizations, and $m=100$ masks and timesteps per $s=1$ full spatial field realizations. The validation data has the following parameters: $\rho=0.05$, $p=5$ parameters from $\tilde{\Theta}$, $s=5$ full spatial field realizations, and $m=1$ masks and timesteps. There are $10$ epochs per data draw and $10$ total data draws. The batch size is 512.
Proportion U-Net
The parameters are fixed as follows: the variance is $\nu=1.5$ and the length scale is $\ell=3$, whereas the proportion of observed locations $\rho\in[0.01,0.5]$ varies. We follow the process in Section S2.2 for simulating data. The training data have the following parameters: $r=50$ observed proportion $\rho$ uniformly sampled from $[0.01,0.525]$, $p=1$ parameters from $\Theta=\tilde{\Theta}=\{3\}\times\{1.5\}$, $s=25$ full spatial field realizations, and $m=100$ masks and timesteps per $s=25$ full spatial field realizations. The validation data have the following parameters: the same $r=50$ observed proportions from the training data, $p=1$ parameters from $\Theta=\tilde{\Theta}=\{3\}\times\{1.5\}$, $s=1$ full spatial field realizations, and $m=1$ masks and timesteps per $s=1$ full spatial field realizations. There are $10$ epochs per data draw and $40$ total draws. The batch size is 2048.
S3.2 Validation
Visualizations
In Figure S1, the simulations from the true and predictive distributions do not match exactly, as expected; however, they display similar spatial patterns and seamlessly incorporate the partially observed field as intended. Additionally, amortization is highly effective. As length scale increases in the left subfigure, the spatial correlations increase in the NCS simulations mirroring the true conditional simulations. No matter the number of observed spatial locations, NCS seamlessly integrates the simulated unobserved values with the partially observed field as shown in the right subfigure. From these observations, we conclude that NCS simulations are visually reasonable as one cannot discriminate between the truth and NCS by the naked eye.
Figure S1: Left panel: NCS with a Gaussian process with variance $\nu=1.5$, length scale $\ell\in\{1,2,3,4,5\}$ (from top to bottom), and proportion of observed locations $\rho=0.05$. Right panel: NCS with a Gaussian process with variance $\nu=1.5$, length scale $\ell=3$, and observed proportions $\rho\in\{0.01,0.05,0.1,0.25,0.5\}$ (from top to bottom). For each panel: Left columns: Observations. Middle columns: Underlying spatial fields (generated using unconditional simulation, only partially observed). Right columns: Conditional simulations with NCS implemented with a parameter U-Net (left panel) or a proportion U-Net (right panel).
Conditional Marginal and Bivariate Densities
The true predictive distributions of a Gaussian process are fully determined by its first and second moments. As such, if the true and NCS conditional univariate and bivariate densities are equivalent up to some approximation error due to finite sample estimation, then the NCS simulations are representative of the true predictive distributions. In Figure S2, the true and NCS conditional univariate and bivariate densities appear equivalent up to approximation error. While Figure S2 only shows densities for a select number of spatial locations and partially observed spatial fields, we compared conditional univariate and bivariate densities for a multitude of spatial locations and observed similar behavior. Consequently, we conclude that the first and second moments of the NCS and true predictive distributions are approximately equivalent, and thus, the NCS simulations are representative of the true predictive distributions.
Figure S2: Left panel: Predictive densities of a Gaussian process with variance $\nu=1.5$, length scale $\ell\in\{1,2,3,4,5\}$ (from top to bottom), and observed proportion $\rho=0.05$. Right panel: As in left panel, but with $\nu=3$, and with observed proportions $\rho\in\{0.01,0.05,0.1,0.25,0.5\}$ (from top to bottom). For each panel: Left columns: Observations and prediction locations for which univariate densities (at the red crosses) and bivariate densities (across the red stars) are visualized. Middle columns: Exact (blue) and NCS (orange) empirical conditional univariate densities at the locations marked with crosses. Right columns: Exact (blue) and NCS (orange) empirical conditional bivariate densities across the locations marked with stars.
Conditional Mean and Correlation Metrics
Figure S3 and Figure S4 contain the true and NCS conditional mean fields and correlation maps which serve as additional metrics to examine whether, and if so, how the first and second moments of the NCS and true predictive distributions differ. In both figures, there are little to no visual differences between the true and NCS empirical conditional mean fields and correlation heatmaps. For both the NCS and true predictive distributions in Figure S4, spatial correlations generally increase with the length scale, unless there are observed locations (marked in white) near the spatial location of interest (marked in yellow). This behavior is best illustrated in the fourth column of the left subfigure; the conditional correlations sharply decrease beyond the observed location (in white) two pixels above the spatial location of interest (in yellow) yet are unaffected below the spatial location of interest due to the absence of observed data in that area. Such behavior is expected for the true predictive distributions, yet is striking for NCS nonetheless and signals that NCS is correctly conditioning on the given partially observed spatial fields.
Figure S3: Left panel: NCS with a Gaussian process with variance $\nu=1.5$, length scale $\ell\in\{1,2,3,4,5\}$} (from top to bottom), and observed proportions $\rho=0.05$. Right panel: NCS with a Gaussian process with variance $\nu=1.5$, length scale $\ell=3$, and observed proportions $\rho\in\{0.01,0.05,0.1,0.25,0.5\}$} (from top to bottom). For each panel: Left columns: Observations. Middle columns: True empirical conditional mean field. Right columns: empirical NCS conditional mean field implemented with a parameter U-Net (left panel) or a proportion U-Net (right panel).
Figure S4: True (left) and NCS (right) empirical conditional correlation heatmaps for spatial location of interest (in yellow) for a Gaussian process with $\nu=1.5$ and $\ell=1,2,3,4,5$ (top to bottom, left subfigure) and $\ell=3$ and observed proportions $\rho=.01,.05,.1,.25,.5$ (top to bottom, right subfigure).
S4 Brown–Resnick: Intractable Case Study
S4.1 Details on Standard Approximation for Brown–Resnick
Further Description
For both FCS and LCS, we utilize the function condrmaxstab in the SpatialExtremes package to approximate the predictive distributions (Ribatet, 2020). This function involves a three-step procedure in which the first step involves identifying feasible partitions of the $k$ observed locations based on whether they share common extremal functions (Dombry et al., 2013). The extremal functions $\varphi_{\bm{s}_{i}}^{+}$ of a max-stable realization $\bm{z}$ are functions $\eta_{i}W_{i}(\bm{s})$ from the stochastic representation of de Haan (1985) in (S7) over the set of spatial locations $\mathcal{S}$ for which the realization $\bm{z}$ and the extremal function $\varphi_{\bm{s}_{i}}^{+}$ evaluated at location $\bm{s}_{i}$ are equal. Extremal functions are not necessarily unique for all spatial locations $\bm{s}_{i}\in\mathcal{S}$ because an extremal function $\varphi_{\bm{s}_{i}}^{+}$ evaluated at another location $\bm{s}_{j}$ may equal the realization $\bm{z}$ at location $\bm{s}_{j}$ as well. As such, this first step requires an approximation via a MCMC algorithm, specifically a Gibbs sampler, to achieve computational tractability because the number of all possible partitions for $k$ locations is the Bell number $B_{k}$ which grows more than exponentially with respect to $k$ (Castruccio et al., 2016). The second and third steps involve simulating extremal and sub-extremal functions respectively from tractable, closed-form distributions and as such, are not computationally intensive. See Dombry et al. (2013) for full mathematical details and Huser et al. (2019) for another application and assessment of the above Gibbs sampler.
As shown in Figure S5, FCS suffers from negative bias for unobserved locations farther from the spatial domain center when the range parameter is small. This is probably due to the insufficient number of subextremal functions drawn by the condrmaxstab function, which affects mostly locations far away from the observed locations when the variogram $\gamma(h)$ increases to infinity.
Implementation
To implement FCS and LCS, we use the condrmaxstab function with the given set of observed locations (FCS) or the nearest seven neighbors to the select unobserved location or locations (LCS), the type of spatial process set to a Brown–Resnick process, the nugget set to .00001, the parameters set to the true parameters, the burn-in period set to 1000, and the thinning factor set to 100. Here, a nugget of .00001 adds an insignificant amount of noise to the Brown–Resnick unconditional simulations used during the third step to address computational problems which occur in the absence of a nugget.
We use univariate and bivariate LCS to produce conditional univariate and bivariate densities and conditional mean fields. Univariate or bivariate LCS refers to conditional simulation for a single select unobserved location or two select unobserved locations given the seven nearest observed neighbors. In the bivariate case, we first select the three nearest neighbors $\mathcal{N}_{3,1},\mathcal{N}_{3,2}$ for each of the two locations, then randomly sample $7-\lvert\mathcal{N}_{3,1}\cup\mathcal{N}_{3,2}\rvert$ locations from the remaining unique seven nearest neighbors for each of the two select locations. This process ensures there are neighbors for each of the two locations of interest in the conditioning set for bivariate LCS.
S4.2 Training Details
The major difference between the Gaussian and Brown–Resnick case studies is the log-transformation of the training and validation data simulated from the Brown–Resnick process. As such, NCS produces conditional simulations on the (log-transformation) Gumbel scale, and all results are on the Gumbel scale.
Hyper-parameters
As in the Gaussian process case study, we did not discover any training sensitivities to any hyperparameters except with respect to amortization—the amortized variable ought to be sampled from a continuous, positive distribution. As such, we used the same hyperparameters as in the Gaussian process case study with some noted exceptions due to computational efficiency.
Parameter U-Net
The details for simulating training and validation data are essentially the same as for the Parameter U-Net in the Gaussian process case study except the number of proportion values is $r=200$ for the training data and the number of full spatial fields is $s=10$ for the validation data. The batch size, number of data draws, and epochs per data draw are also the same.
Proportion U-Net
The details for simulating training and validation data are essentially the same as for the Proportion U-Net in the Gaussian process case study except the full spatial fields is $s=25$ for the validation data. The batch size, number of data draws, and epochs per data draw are also the same.
Small Conditioning Set U-Net
We train one U-Net for each range parameter $\lambda\in\{1,2,3,4,5\}$}. However, the smoothness is fixed at $\nu=1.5$ and the minimum and maximum number of observations is fixed at one and ten respectively for all U-Nets. For evaluation, the number of observed locations does not exceed seven because FCS as implemented in the SpatialExtremes R package is fragile beyond seven. To simulate training and validation data, we follow the process described in Section S2.2. The training data has the following parameters: the minimum and maximum number of observations $o_{L}=1$ and $o_{U}=10$, $s=32$ full spatial field realizations, and $m=100$ conditioning sets, masks and timesteps per $s=32$ full spatial field realizations. The validation data has the following parameters: the minimum and maximum number of observations $o_{L}=3$ and $o_{U}=3$, $s=32$ full spatial field realizations, and $m=1$ conditioning sets, masks and timesteps per $s=32$ full spatial field realizations. There are $10$ epochs per data draw and $40$ total data draws. The batch size is $2048$.
S4.3 Validation
@@
Figure S5: True fully observed (left) fields and FCS (middle) and NCS (right) simulations of a Brown–Resnick Process with smoothness $\nu=1.5$ and range $\ell=1,2,3,4,5$ (top to bottom) on the Gumbel scale for three observed locations outlined in red. Values outside [2,6][-2,6] have been replaced with the closest value in this interval.
In Figure S5, both FCS (middle) and NCS (bottom) simulations are shown because the number of observed locations is less than eight. Clearly, FCS suffers from negative bias for unobserved locations farther from the spatial domain center when the range parameter is small. This is due to the way MCMC is implemented for this approximation method. Due to this negative bias, the NCS simulations appear more accurate overall, although NCS suffers from slightly increased graininess as the range parameter increases.
@@
Figure S6: True partially observed spatial fields (left) and NCS (right) empirical conditional mean fields of a Brown–Resnick process with smoothness $\nu=1.5$ and range $\lambda=1,2,3,4,5$ (top to bottom) and observed proportions $\rho=.05$ (left subfigure) and $\lambda=3$ and observed proportions $\rho=.01,.05,.1,.25,.5$ (top to bottom, right subfigure) on the Gumbel scale. Values outside [2,6][-2,6] have been replaced with the closest value in this interval.
In Figure S6, there are no discernible differences between the univariate LCS and NCS empirical conditional mean fields.

### v2

@@ Acknowledgments
Fix parameter $\bm{\theta}=(\lambda,\nu)^{\prime}$.
Select a minimum and maximum number of observed locations, $o_{L}<o_{U}\in\mathbb{N}^{+}$.
for $o_{j}$ in $o_{L},\ldots,o_{U}$ do
 Simulate unconditionally $\{\bm{x}_{0,k}\}_{k\in[s]}\sim p(\cdot\mid\bm{\theta})$
@@
  $\mathring{\mathcal{S}}_{l}=\{\bm{s}_{l_{i}}\in\mathcal{S}\mid\bm{s}_{l_{i}}\sim\textrm{U}(\mathcal{S}),i\in[o_{j}]\}$
  $t_{l}\sim U(\{0,1,\dots,T\})\textrm{ for }l\in[m]$
  $\bm{\mathring{x}}_{0,l}=(x_{0,k,j}:M_{j}(\mathring{\mathcal{S}}_{l})=1)^{\prime}\textrm{ for }l\in[m]$
  $\tilde{\bm{x}}_{0,l}=(x_{0,k,j}:M_{j}(\mathring{\mathcal{S}}_{l})=0)^{\prime}\textrm{ for }l\in[m]$
  $\tilde{\bm{x}}_{t_{l}}\sim p(\cdot\mid\bm{\mathring{x}}_{0,l},\tilde{\bm{x}}_{0,l}):=\mathcal{N}(\sqrt{\underline{\alpha}_{t}}\tilde{\bm{x}}_{0,l},\underline{\sigma}_{t}^{2}\tilde{\bm{I}})\textrm{ for }l\in[m]$
  $\mathcal{D}\leftarrow\mathcal{D}\cup\{\{({\bm{\mathring{x}}_{0,l}}^{\prime},{\tilde{\bm{x}}_{t_{l}}}^{\prime})^{\prime},\bm{M}(\mathring{\mathcal{S}}_{l}),\bm{\theta}_{j},t_{l}\}:l\in[m]\}$
@@
Specifically, we train one U-Net for each specific range parameter $\lambda\in\{1,2,3,4,5\}$}. However, we fix the smoothness to $\nu=1.5$ and the minimum and maximum number of observations to one and ten respectively for all U-Nets. For validation, the number of observed locations does not exceed seven because FCS as implemented in the SpatialExtremes R package is fragile beyond seven. To simulate training and validation data, we follow the process described in Section S2.2. For simulating training data, the minimum and maximum number of observations were set to $o_{L}=1$ and $o_{U}=10$, $s=32$ full spatial field realizations were generated per number of observations between $o_{L}$ and $o_{U}$, and $m=100$ conditioning sets, masks and timesteps were simulated for each full spatial field realization.
S4.3 Validation Details
Conditional Validation Dataset (Small Conditioning Set)
@@
For each range value $\lambda\in\{1,2,3,4,5\}$ and each conditioning set size $m$ ranging from one to seven, we unconditionally simulated a reference spatial field $\bm{x}_{0}$ from a Brown–Resnick process with the given range $\lambda$ and fixed smoothness $\nu=1.5$. For each reference spatial field, we generated a conditioning set $\mathring{\mathcal{S}}$ and corresponding mask $\bm{M}(\mathring{\mathcal{S}})$ by uniformly sampling from $\mathcal{S}$ without replacement, such that $\lvert\mathring{\mathcal{S}}\rvert=m$. Using each of the $35$ partially observed fields $\bm{\mathring{x}}_{0}=(x_{0,j}:M_{j}(\mathring{\mathcal{S}})=1)^{\prime}$, we generated $m=4000$ full spatial field replicates $\{(\bm{\mathring{x}}_{0}^{\prime},\tilde{\bm{x}}_{0,i}^{\prime})^{\prime}\}_{i\in[m]}$ with NCS and FCS (an MCMC approximation described in Section 3.1). We use this dataset in Section 3.2 to validate NCS when the number of observed locations does not exceed seven.
Unconditional Validation Dataset (Small Conditioning Set U-Net)
For the validation method described in Section 2.6, we need to generate unconditional samples from the conditional ones. As discussed in Section 2.6, we do this by generating unconditionally at observed locations and then conditionally simulating at unobserved locations; the resulting concatenated sample is then one from the unconditional distribution. Algorithm S3 summarizes how we do this, using $m=4000$ NCS-approximated or FCS-approximated unconditional simulations using varying range $\lambda\in\{1,2,3,4,5\}$}, fixed smoothness $\nu=1.5$, and conditioning sets with varying sizes ranging from one to seven.
@@
Algorithm S3 Generating NCS-approximated Unconditional Validation Dataset (Small Conditioning set)
@@
Fix parameter $\bm{\theta}=(\lambda,\nu)^{\prime}$ and number of observed locations $o_{j}$ and set $\mathcal{D}_{\textrm{val}}=\emptyset$
for $i$ in $1,\dots,m$ do
@@
$\mathring{\mathcal{S}}_{i}\sim\textrm{U}(\bm{\mathcal{S}}_{o_{j}})\textrm{ where }\bm{\mathcal{S}}_{o_{j}}=\{\mathcal{S}_{i}\subset\mathcal{S}:\lvert\mathcal{S}_{i}\rvert=o_{j}\}$
$\tilde{\bm{x}}_{0,i}\sim\tilde{p}_{0}(\cdot\mid\bm{\mathring{x}}_{0,i},\bm{\theta})\textrm{ for }\bm{\mathring{x}}_{0,i}=(x_{0,i,j}:M_{j}(\mathring{\mathcal{S}}_{i})=1)^{\prime}\textrm{ and NCS/FCS distr.\ }\tilde{p}_{0}(\cdot\mid\bm{\mathring{x}}_{0,i},\bm{\theta})$
$\mathcal{D}_{\textrm{val}}\leftarrow\mathcal{D}_{\textrm{val}}\cup\{(\bm{\mathring{x}}_{0,i}^{\prime},\tilde{\bm{x}}_{0,i}^{\prime})^{\prime}\}$
@@
Return $\mathcal{D}_{\textrm{val}}=\{(\bm{\mathring{x}}_{0,i}^{\prime},\tilde{\bm{x}}_{0,i}^{\prime})^{\prime}\}_{i\in[m]}$
Unconditional Validation Dataset (Parameter and Proportion U-Net)
We apply the process described in Section 2.6 and the previous paragraph but with varying parameters and observed proportions as shown in Algorithm S4. We generate $m=4000$ NCS-approximated unconditional simulations using either varying range $\lambda\in\{1,2,3,4,5\}$, fixed smoothness $\nu=1.5$, and fixed observed proportion $\rho=.05$ for the parameter U-Net, or varying observed proportion $\rho\in\{0.01,0.05,0.1,0.25,0.5\}$ with fixed range $\lambda=3$ and smoothness $\nu=1.5$ for the proportion U-Net.
Algorithm S4 Generating NCS-approximated Unconditional Validation Dataset (Parameter and Proportion U-Nets)
Fix parameter $\bm{\theta}=(\lambda,\nu)^{\prime}$ and observed proportion $\rho$ and set $\mathcal{D}_{\textrm{val}}=\emptyset$
Set the number of NCS-approximated unconditional replicates $m$.
for $i$ in $1,\dots,m$ do
@@
$\bm{M}(\mathring{\mathcal{S}}_{i})\sim\textrm{Bin}(n,\rho)$ with $n=32^{2}$
$\tilde{\bm{x}}_{0,i}\sim\tilde{p}_{0}(\cdot\mid\bm{\mathring{x}}_{0,i},\bm{\theta})\textrm{ for }\bm{\mathring{x}}_{0,i}=(x_{0i,j}:M_{j}(\mathring{\mathcal{S}}_{i})=1)^{\prime}$ and NCS/LCS distr. $\tilde{p}_{0}(\cdot\mid\bm{\mathring{x}}_{0,i},\bm{\theta})$
$\mathcal{D}_{\textrm{val}}\leftarrow\mathcal{D}_{\textrm{val}}\cup\{(\bm{\mathring{x}}_{0,i}^{\prime},\tilde{\bm{x}}_{0,i}^{\prime})^{\prime}\}$
@@
Return $\mathcal{D}_{\textrm{val}}=\{(\bm{\mathring{x}}_{0,i}^{\prime},\tilde{\bm{x}}_{0,i}^{\prime})^{\prime}\}_{i\in[m]}$
@@
The metrics in Section S2.3.2 such as the conditional univariate and bivariate densities, conditional mean fields, and conditional correlation metrics only compare low-dimensional properties of NCS to those of the true predictive distributions or standard approximations. Here, we discuss metrics that implicitly validate the high-dimensional patterns of NCS-approximated predictive distributions via NCS-approximated unconditional simulations. These metrics, referred to as multivariate metrics, are evaluated using the aforementioned unconditional validation datasets. For the Brown–Resnick case study, where these multivariate metrics are especially relevant, we considered the extremal correlation function and the spatial minimum, maximum, and absolute summation in any given full realization. Distributions of these quantities are shown in our results (e.g., Figure 5) for different parameters and missingness cases.
S4.4 Results
@@
Figure S5: True fully observed (left) fields and FCS (middle) and NCS (right) simulations of a Brown–Resnick Process with smoothness $\nu=1.5$ and range $\ell=1,2,3,4,5$ (top to bottom) on the Gumbel scale for three observed locations shown in red. Values outside [2,6][-2,6] have been replaced with the closest value in this interval.
In Figure S5, we show results for both FCS (middle) and NCS (right) simulations for the case when three locations are observed. Clearly, FCS suffers from negative bias for unobserved locations far away from the spatial domain center when the range parameter is small. This is due to the way MCMC is implemented for this approximation method. Due to this negative bias, the NCS simulations appear more representative overall, although NCS suffers from slightly increased graininess as the range parameter increases.
Figure S6: True partially observed spatial fields (left) and NCS (right) empirical conditional mean fields of a Brown–Resnick process with smoothness $\nu=1.5$ and range $\lambda=1,2,3,4,5$ (top to bottom) and observed proportions $\rho=.05$ (left panel) and $\lambda=3$ and observed proportions $\rho=.01,.05,.1,.25,.5$ (top to bottom, right panel) on the Gumbel scale. Values outside [2,6][-2,6] have been replaced with the closest value in this interval.
@@
In Figure S6, we show results for various spatial process parameter and proportion settings. We see no discernible differences between the univariate LCS and NCS empirical conditional mean fields.